A deep-sea diver uses a gas cylinder with a volume of 10.0 L and a content of 51.0 g of O2 and 33.2 g of He You may want to reference (山-pages 410-412) section 10.6 while completing this problem. Part A Calculate the partial pressure of each gas and the total pressure if the temperature of the gas is 18 oC. Express the pressures in atmospheres to three significant digits separated by commas ANSWER: atm
Answer #1

Molar mass of O2 = 32 g/mol

51 g / 32 g/mol => 1.59 moles

Molar mass of He = 4 g/mol

33.2 g / 4 g/mol => 8.3 moles

Total moles = 1.59 + 8.3 => 9.89 moles

Total P = nRT / V

P = 9.89 moles x 0.082 L atm / mol.K x 291 K / 10 L ( note 18 C = 18 + 273 = 291 K )

P = 23.60 atm

partial P of O2 => (1.59 / 9.89 ) x 23.6 => 3.79 atm

partial P of He => (8.3 / 9.89 ) x 23.6 => 19.80 atm

Answer :  3.79, 19.8, 23.6